Below are the ages of 20 winners of the Acadeiny Award for Best Actor and Best Actress. Best Actor: 42 54 52 37 38 60 32 45 60 32 45 60 46 40 36 50 47 29 43 37 Best Actress: 32 4 1 33 31 74 33 32 22 49 61 21 41 26 33 44 80 42 29 33 4) Create a back to back stem and leaf plot. TABLE 2. Back to back stem and leaf for best actor/actress Best Actor Best Actress 5) Compare the two sets of data, what can be concluded from the stem and leaf display?
A sample of 521 adults was asked How often do you dine out?, the results of the survey are listed in table 1. TABLE 1. How often do you dine out? Response Several times a week Once or twice a week A few times a month Very rarely Never Prequency 103 204 130 79 1) What proportion of adults in the study very rarely dine out? 2) If someone said 20% of the people survey said they dine out several times a week is this a discrpitive statement or an inferential state- ment? (circle one) a) Descriptive b) Inferential 3) In your own words why is using relative frquency better to com- pare different surveys?
Answer #1

1) Proportion of adults that very rarely dine out= No. of adults(frequency) that very rarely dine out /Total no. of adults

= 79/521 = 0.1516

In percentage terms, it is 0.1516*100= 15.16%

2) a) Descriptive.

If someone says the quote as mentioned in the question, it is a Descriptive statement as it can be verified from the data of the sample that has been provided(and can be proofed as below):

Proportion of adults that dine out several times a week =

No. of adults(frequency) that dine out several times a week /Total no. of adults

= 103/521 =0.1976

In percentage terms, it is 0.1976*100= 19.76% = 20% approx

3) Relative frequency is better to compare different surveys because it helps us in understanding the association between variables in a better and a meaningful way. Relative frequency shows percentages/proportions rather than the absolute counts which can be used for direct comparison.

Hence, they help in summarising the data better.

4) Back to back stem and leaf plot

TABLE 2
BEST ACTOR BEST ACTRESS
9 2 1 2 6 9   
8 7 7 6 2 2 3 1 2 2 3 3 3 3 8
7 6 5 5 3 2 0 4 1 1 2 4 9
4 2 0 5
0 0 0 6 1
7 4
8 0
9